An unknown radioactive material has an absorbed dose of 500 μ Gy and gives a equivalent dose of 1 mSv. Calculate the radiation weighting factor of the material.



1
Expert's answer
2021-03-08T13:32:05-0500

Equivalent dose "H" is calculated using the mean absorbed dose "D"  multiplied by the radiation weighting factor "W_R"  which is dependent on the type and energy of the radiation:


"H = W_RD"

expressing the radiation weightting factor and substituting "D = 500\mu Gy = 5\times 10^{-4}Gy", and "H = 1mSv = 1\times 10^{-3}Sv", obtain:


"W_R = \dfrac{H}{D}\\
W_R = \dfrac{1\times 10^{-3}}{5\times 10^{-4}} = 2"

Answer. 2.